# Quantitative analysis of the value of investment in research facilities, with examples from cyberinfrastructure

### v2

@@ Discussion
The core, long-standing, and common result of “lease vs. buy” analyses is that it is financially more efficient to operate on-premises research computing systems (or, as noted earlier, to house university-owned systems in owned or leased remote data centers). That is, ROIproxy for on-premises operations is well over 1. For example, [10] found in 2018 that the cost to outsource to a cloud provider would have been between 2.5 and 3.7 times greater than the cost of owning and operating an on-premises supercomputer. They found roughly similar results when comparing outsourcing with an on-premise cloud system - outsourcing being 1.9 to 3.1 times more expensive than an on-premise cloud. Several analyses of the cost-effectiveness of XSEDE now consistently show a calculated ROIproxy well over 1. That is, it is far less expensive for the USA federal government to offer grants and contracts to universities to provide cyberinfrastructure services than it would be to purchase such services on the open market within the USA. These ROI figures are somewhat higher than the figure of 1.8 found in an analysis of a recent upgrade to CERN facilities [22]. However, considering the differences in methodology, this must be considered a reasonable consistency among these findings.
@@
Looking at Purdue University specifically, the analysis presented by Smith offers clear guidance regarding the benefits of investing in equipment and staff. Today, USD 100,000 will pay for compute node servers providing approximately 200 TeraFLOPS of FP64 Tensor Core (a specific nVidia product described at [43]) performance. A leader at Purdue University could conclude that an investment of USD 100,000 in computational capability will correspond with an increase in total university research expenditures by the institution of USD 6.45M. This implies a significant increase in receipts of extramural grant funding. Similarly, USD 100,000 is roughly the full annual cost of one entry-level RCD professional (salary and benefits). Every USD 100,000 invested in staff costs – adding one person – corresponds to an increase of USD 14M in total grant expenditures. Indeed, at Purdue University, which is strongly focused on engineering, the analysis by Smith suggests that as of the time of his analysis, access to computing power was an enabler and a limiting factor of research activity at a very large scale within the university community. The fact that relationships between academic and financial output metrics and independent variables related to investment in computational facilities and RCD staff remain statistically significant when analyzed across six different R1 universities is strong testament to the reality of an underlying causal relationship. Of course, what will be meaningful is the relationship within a particular university. For higher education leaders trying to react sensibly and strategically to budget challenges, the production-function approach seems to have much to offer.
@@
There are relatively few peer-reviewed analyses of the academic uses of commercial cloud services as a research-enabling facility. (Publications that give the results of surveys by cloud vendors and so-called industry observers are easy to find, but they are, typically, essentially advertisements, not scholarly research.) Three recent systematic and peer-reviewed surveys of commercial cloud use are [44], [45], and [46]. The Coalition for Academic Scientific Computation (CASC) is an organization focused on the research uses of cyberinfrastructure in higher education in the USA [47]. The two CASC surveys focus on use of commercial clouds in supporting research activities. These surveys have confirmed a rise over time in the academic use of cloud facilities in support of research activities. The 2021 survey, in particular, found that the reasons for such use were primarily related to functional issues rather than cost. In that survey, the most frequent reasons given for the purchase of commercial cloud resources in support of research activities were (in decreasing order of frequency, out of a total of 69 responses):
@@
Availability of public funding is a key enabler of publicly funded research, and as such there has historically been much less such research in the Global South (cf. [48]). Indeed, in some areas there has been considerable progress in publication of and citations to research led by researchers in the Global South (for example, as regards sustainability research, see [49]). There have been notable publications taking a return on investment approach to evaluating medical interventions in the global south. For example, [50] analyzed the cost-effectiveness of surgical interventions in patient care in a conflict-affected area of South Sudan and found a ROI of 14. ROI for healthcare intervention in the form of deploying health extension workers in Ethiopia found an ROI of 1.3 to 3.7 [51]. We have not found any ROI analyses for research activities in the Global South, but clearly this concept is well understood and used. The quantitative methods described in this paper were a good deal of work to develop and deploy in the form of first proofs of concept. Now that this has been done, applying them will we hope be more straightforward. It is the hope of the authors that the methods presented here will be of use to researchers and research administrators in the Global South as they seek to justify additional funding for research about the Global South, led by researchers from the Global South.
As regards the Global North, there are many reasons to hope that the current state of uncertainty about funding for university-based research comes to an end. By the same token, there are many reasons to hope that the wave of higher ed closures seen in the USA is not replicated in the EU. We hope that the presentation of methods for quantitatively expressing the value of university-based research institutions - straightforwardly extensible to university-based research generally, will become useful tools in addressing current questions about the value of such research. A reasonable critic might offer the observation that some perhaps significant portion of today’s skepticism regarding higher education comes from a failure in the past to analyze cost-benefit ratios according to various different perspectives and value systems. Indeed, in the USA, it was possible in the 1980s to justify the then-new NSF supercomputer program entirely based on the potential new knowledge it would create. The creation of new knowledge, not possible to create without use of a supercomputer, was considered sufficient without any reference to cost/effectiveness at all [52] that was impossible to obtain without use of a supercomputer. The methods demonstrated here might, if widely applied, lead to greater consensus regarding at least this aspect of higher education institutions in the Global North.
@@
Fig 1 presents an example of Balanced Scorecard template for the assessment of research-enabling facilities. This example is made as general as possible for made as general as possible. This depiction shows rows of metrics applicable to any sort of research-enabling facilities where general exemplar metrics can be suggested. In places where metrics have to be more specific in order to make sense, for example in operational metrics, examples from cyberinfrastructure facilities are shown. Such operational metrics can be altered to express the essential measures for any kind of research-enabling facilities. Following Kaplan and Norton’s discussion, all metrics should somehow be related to organizational strategies and tactics and may include statements of quantitative goals set (e.g., key performance indicators). This example, with little modification, should be useful to anyone trying to argue for the value of any cyberinfrastructure facility supporting research.
@@
Fig 1: Balanced Scorecard for research-enabling facilities. Metrics applicable to any sort of research-enabling facilities are shown in plain text. Italics are used for operational metrics which are by their nature specific to a particular sort of facility. For purposes of this example, metrics applicable to a cyberinfrastructure facility are shown. The arrows are as shown in Kaplan and Norton’s orignal example and indicate the interrelatedness of organizational performance in the four general categories they specify. Particular notes regarding entries in this table: 1It is generally possible to estimate total societal value only after a facility has been in operation for some time. 2 The operational metrics shown here are specific to cyberinfrastructure facilities. There will be discipline-specific operational metrics relevant to the mission of any research-enabling facility.
@@
As long as finances remain a critical concern in higher education - and there is no end in sight to concerns about finances in higher education - financially based assessments are likely to be of value in practical decision-making and in conveying the value of investments in research facilities and research generally to the taxpaying public. It seems highly likely that it will be valuable to continue extending the methods of evaluation.
@@ Conclusions
Credible hypotheses about the linkage between economic growth and investment in innovation have existed at least since Schumpeter’s 1912 Theorie der wirtschaftlichen Entwicklung (Theory of Economic Development). Recent analyses qualitatively suggest a relationship between investment in information technology and economic benefits at a national level [18]. It is far more difficult to assess the value of any particular instance of a research-enabling facility, such as a cyberinfrastructure facility. However, doing so is essential for setting policy and investment plans for any particular type of infrastructure at any given institution or within any given national research sector. Even more essential today may be using such tools to demonstrate the financial value of such investments to potentially skeptical stakeholders, including elected officials and taxpayers.
We have presented here several approaches to assessing and conveying the value of investment in particular facilities. We can summarize these approaches as follows:
@@
  • Production-function analyses, which can demonstrate relationships between investments in facilities and academic and financial outputs and which are particularly valuable in policy setting and management within particular organizations.
All of these methods may be applied to any sort of research-enabling or research-facilitating infrastructure. Some of these methods are of interest only within academia but stand to be highly valuable, within those academic institutions, for decision-making about research-enabling facility investments. Many of the above methods address the financial value of investment in research facilities at the regional or national level. These methods may be valuable when attempting to convey, to stakeholders ranging from individual voters to elected officials, the value - to society, to the economy, or to the global competitiveness of a nation as a whole - of investment in research. Most of these methods may be applied retroactively, given the existence of sufficiently good recordkeeping. These approaches, combined with the Balanced Scorecard as a way to depict outcomes, offer important tools for research policy planning and analysis. These tools also offer the potential to make convincing a case for the importance of public investment in research and research-enabling facilities even to constituents and stakeholders who will consider only returns states in financial terms as significant. As institutions of higher education in the USA and EU face important and difficult decisions regarding allocation of funds internally, and skeptical stakeholders externally, clarity in demonstrating value is important. In the global south, research institutions and institutions of higher education dedicated to research programs that further advances in quality of life and in the economy may find these techniques applicable in arguing for expanded investment in research. Stakeholder support for government funding of research depends on many factors, of which one essential factor is clarity in expressing the value of research carried out by institutions of higher education.
@@ Acknowledgments
We thank Kristol Hancock, Marie Deer, and Tracey Metivier for editing. Thanks to Katja Bookwalter, who created Figure 1. The research presented here was supported in part by grant awards from the USA National Science Foundation (awards #1548562, #2227627, #1445604, #2005506, #2005632) and has also been supported in part by the Indiana University Pervasive Technology Institute. Any opinions expressed here are those of the authors and may not necessarily represent the views of any funding organizations.
@@
Data availability statement: No new data presented. Author contributions (CRediT): Conceptualization - all; Formal analysis - all; Funding acquisition - Snapp-Childs, Smith, Towns; Investigation - all; Writing original draft, review, editing - all. Generative AI statement: No generative AI used.

### v3

@@ Discussion
The core, long-standing, and common result of “lease vs. buy” analyses is that it is financially more efficient to operate on-premises research computing systems (or, as noted earlier, to house university-owned systems in owned or leased remote data centers). That is, ROIproxy for on-premises operations is well over 1. For example, [10] found in 2018 that the cost to outsource to a cloud provider would have been between 2.5 and 3.7 times greater than the cost of owning and operating an on-premises supercomputer. They found roughly similar results when comparing outsourcing with an on-premises cloud system - outsourcing being 1.9 to 3.1 times more expensive than an on-premises cloud. Several analyses of the cost-effectiveness of XSEDE now consistently show a calculated ROIproxy well over 1. That is, it is far less expensive for the USA federal government to offer grants and contracts to universities to provide cyberinfrastructure services than it would be to purchase such services on the open market within the USA. These ROI figures are somewhat higher than the figure of 1.8 found in an analysis of a recent upgrade to CERN facilities [22]. However, considering the differences in methodology, this must be considered a reasonable consistency among these findings.
@@
Looking at Purdue University specifically, the analysis presented by Smith offers clear guidance regarding the benefits of investing in equipment and staff. Today, USD 100,000 will pay for compute node servers providing approximately 200 TeraFLOPS of FP64 Tensor Core (a specific nVidia product described at [44]) performance. A leader at Purdue University could conclude that an investment of USD 100,000 in computational capability will correspond with an increase in total university research expenditures by the institution of USD 6.45M. This implies a significant increase in receipts of extramural grant funding. Similarly, USD 100,000 is roughly the full annual cost of one entry-level RCD professional (salary and benefits). Every USD 100,000 invested in staff costs – adding one person – corresponds to an increase of USD 14M in total grant expenditures. Indeed, at Purdue University, which is strongly focused on engineering, the analysis by Smith suggests that as of the time of his analysis, access to computing power was an enabler and a limiting factor of research activity at a very large scale within the university community. The fact that relationships between academic and financial output metrics and independent variables related to investment in computational facilities and RCD staff remain statistically significant when analyzed across six different R1 universities is a strong testament to the reality of an underlying causal relationship. Of course, what will be meaningful is the relationship within a particular university. For higher education leaders trying to react sensibly and strategically to budget challenges, the production-function approach seems to have much to offer.
@@
There are relatively few peer-reviewed analyses of the academic uses of commercial cloud services as a research-enabling facility. (Publications that give the results of surveys by cloud vendors and so-called industry observers are easy to find, but they are, typically, essentially advertisements, not scholarly research.) Three recent systematic and peer-reviewed surveys of commercial cloud use are [45], [46], and [47]. The Coalition for Academic Scientific Computation (CASC) is an organization focused on the research uses of cyberinfrastructure in higher education in the USA [48]. The two CASC surveys focus on use of commercial clouds in supporting research activities. These surveys have confirmed a rise over time in the academic use of cloud facilities in support of research activities. The 2021 survey found that the reasons for such use were primarily related to functional issues rather than cost. In that survey, the most frequent reasons given for the purchase of commercial cloud resources in support of research activities were (in decreasing order of frequency, out of a total of 69 responses):
@@
Availability of public funding is a key enabler of publicly funded research, and as such there has historically been much less such research in the Global South (cf. [49]). In fact, in some areas there has been considerable progress in the publication of and citations to research led by researchers in the Global South (for example, in terms of sustainability research, see [50]). There have been notable publications taking a return on investment approach to evaluating medical interventions in the Global South. For example, [51] analyzed the cost-effectiveness of surgical interventions in patient care in a conflict-affected area of South Sudan and found a ROI of 14. ROI for healthcare intervention in the form of deploying health extension workers in Ethiopia found an ROI of 1.3 to 3.7 [52]. We have not found any ROI analyses for research activities in the Global South, but clearly this concept is well understood and used. The quantitative methods described in this paper were a good deal of work to develop and deploy in the form of first proofs of concept. Now that this has been done, applying them will, we hope, be more straightforward. It is the hope of the authors that the methods presented here will be of use to researchers and research administrators in the Global South as they seek to justify additional funding for research about the Global South, led by researchers from the Global South. This hope is not to imply that the Global South is itself homogeneous but there are many problems faced in the Global South where researchers and countries share more in common with each other than with researchers and countries in the Global North.
As regards the Global North, there are many reasons to hope that the current state of uncertainty about funding for university-based research comes to an end. By the same token, there are many reasons to hope that the wave of higher ed closures seen in the USA is not replicated in the EU. We hope that the presentation of methods for quantitatively expressing the value of university-based research institutions - straightforwardly extensible to university-based research generally, will become useful tools in addressing current questions about the value of such research. A reasonable critic might offer the observation that some perhaps significant portion of today’s skepticism regarding higher education comes from a failure in the past to analyze cost-benefit ratios according to different perspectives and value systems. Indeed, in the USA, it was possible in the 1980s to justify the then-new NSF supercomputer program entirely based on the potential new knowledge it would create. The creation of new knowledge, not possible to create without use of a supercomputer, was considered sufficient without any reference to cost/effectiveness at all [53] that was impossible to obtain without use of a supercomputer. The methods demonstrated here might, if widely applied, lead to greater consensus regarding at least this aspect of higher education institutions in the Global North.
@@
Fig 1 presents an example of Balanced Scorecard template for the assessment of research-enabling facilities. This example is as general as possible. This depiction shows rows of metrics applicable to any sort of research-enabling facilities where general exemplar metrics can be suggested. In places where metrics must be more specific in order to make sense, for example in operational metrics, examples from cyberinfrastructure facilities are shown. Such operational metrics can be altered to express the essential measures for any kind of research-enabling facilities. Following Kaplan and Norton’s discussion, all metrics should somehow be related to organizational strategies and tactics and may include statements of quantitative goals set (e.g., key performance indicators). Note that the arrows are as shown in Kaplan and Norton’s original example and indicate the interrelatedness of organizational performance in the four general categories they specify. This example, with little modification, should be useful to anyone trying to argue for the value of any cyberinfrastructure facility that supports research.
@@
Fig 1: Balanced Scorecard for research-enabling facilities. Metrics applicable to any sort of research-enabling facilities are shown in plain text. Italics are used for operational metrics which are by their nature specific to a particular sort of facility. 1It is generally possible to estimate total societal value only after a facility has been in operation for some time. 2 There will be discipline-specific operational metrics relevant to the mission of any research-enabling facility.
@@
As long as finances remain a critical concern in higher education - and there is no end in sight to concerns about finances in higher education - financially based assessments are likely to be of value in practical decision making and in conveying the value of investments in research facilities and research generally to the taxpaying public. It seems highly likely that it will be valuable to continue extending the methods of evaluation.
@@ Conclusions
Credible hypotheses about the linkage between economic growth and investment in innovation have existed at least since Schumpeter’s 1912 Theorie der wirtschaftlichen Entwicklung (Theory of Economic Development). Recent analyses qualitatively suggest a relationship between investment in information technology and economic benefits at a national level [18]. It is far more difficult to assess the value of any particular instance of a research-enabling facility, such as a cyberinfrastructure facility. However, doing so is essential for setting policy and investment plans for any type of infrastructure at any given institution or within any given national research sector. Even more essential today may be using such tools to demonstrate the financial value of such investments to potentially skeptical stakeholders, including elected officials and taxpayers.
We have presented here several approaches to assessing and conveying the value of investment in facilities. We can summarize these approaches as follows:
@@
  • Production-function analyses, which can demonstrate relationships between investments in facilities and academic and financial outputs. Such analyses can be uniquely valuable for policy setting and management within organizations.
These methods may be applied to any type of research-enabling or research-facilitating infrastructure. Some of these methods are of interest only within academia but stand to be highly valuable, within those academic institutions, for decision making about research-enabling facility investments. Many of the above methods address the financial value of investment in research facilities at the regional or national level. These methods may be valuable when attempting to convey, to stakeholders ranging from individual voters to elected officials, the value - to society, to the economy, or to the global competitiveness of a nation as a whole - of investment in research. Most of these methods may be applied retroactively, given the existence of sufficiently good recordkeeping. These approaches, combined with the Balanced Scorecard as a way to depict outcomes, offer important tools for research policy planning and analysis. These tools also offer the potential to make convincing a case for the importance of public investment in research and research-enabling facilities even to constituents and stakeholders who will consider only returns stated in financial terms as significant. As institutions of higher education in the USA and EU face important and difficult decisions regarding allocation of funds internally, and skeptical stakeholders externally, clarity in demonstrating value is important. In the Global South, research institutions and institutions of higher education dedicated to research programs that further advances in quality of life and in the economy may find these techniques applicable in arguing for expanded investment in research. Stakeholder support for government funding of research depends on many factors, of which one essential factor is clarity in expressing the value of research carried out by institutions of higher education.
@@ Acknowledgments
Thanks to Kristol Hancock, Marie Deer, and Tracey Metivier for editing. Thanks to Katja Bookwalter, who created Figure 1. The research presented here was supported in part by grant awards from the USA National Science Foundation (awards #1548562, #2227627, #1445604, #2005506, #2005632) and has also been supported in part by the Indiana University Pervasive Technology Institute. Any opinions expressed are those of the authors and may not necessarily represent the views of any funding organizations.
@@
Data availability statement: No new data presented. Data sources for works cited by coauthors are available from sources cited in those papers/ Author contributions (CRediT): Conceptualization - all; Formal analysis - all; Funding acquisition - Snapp-Childs, Smith, Towns; Investigation - all; Writing original draft, review, editing - all. Generative AI statement: No generative AI used.